The market risk premium is 8%. The market variance is 20%. The variance of asset XYZ is 40%. The covariance between asset XYZ and the market is 10%. According to the CAPM, what should the risk premium of asset XYZ be?
Answer #1

Beta of XYZ = CoVarianceof XYZ And Market Varianceof XYZ

= 0.10 / 0.40 = 0.25

Risk Premium = Beta * market Risk Premium

= 0.25 * 8%

= 2%   

NOTE: The answer to your question has been given below/above. If there is any query regarding the answer, please ask in the comment section. If you find the answer helpful, do upvote. Help us help you.